OpenStudy (anonymous):

What is the sum of a 12-term arithmetic sequence where the last term is 13 and the common difference is –10? 605 660 748 816

OpenStudy (saifoo.khan):

Do you know the sum formula?

OpenStudy (saifoo.khan):

\[S_t = \frac{n}{2} (a+l)\]Where a is the forst term. L is the last. n is the term of total terms.

OpenStudy (saifoo.khan):

i meant, n is the number of terms.

OpenStudy (anonymous):

saifoo in which grade u r now?

OpenStudy (saifoo.khan):

@rockstar111 , heard of A-levels?

OpenStudy (saifoo.khan):

Wait! you can use another formula as well. xD

OpenStudy (anonymous):

A-LEVELS what is this?

OpenStudy (saifoo.khan):

These are two formulas, http://intmstat.com/series-binomial-theorem/Image1702.gif Use second one.

OpenStudy (campbell_st):

ist you need to find the value of the1st term using the last term information, and common difference \[13= a + (12 -1) -10\] \[13 = a- 110\] a = 123 then you can find the sum \[s _{n}= \frac{n}{2}(a + l)\]

OpenStudy (saifoo.khan):

@rockstar111 Let's head to chat and do there.

OpenStudy (anonymous):

i think its 816, am i right?

OpenStudy (campbell_st):

thats it